What is the equation of Five times a number decreased by eight is equal to thirity-two? and also A number multiplied by 8 and divided by four gives 7 more than the number.Thanks!
Answer by mananth(16949)

let number be x
5 times = 5x
less 8
5x-8 = 32
add 8
5x=25
/5
x= 5
---------
Sumilarly
8x/4 =7+x
multiply equation by 4
8x=28+4x
8x-4x=28
4x=28
x=7
m.ananth@hotmail.ca